OpenStudy (anonymous):

help please!! 1 medal and fan:) In TRIANGLE ABC, centroid D is on median AM . AD = x + 5 and DM = 3x – 5. Find AM. A. 6 B. 12 C. 13 D. 4

OpenStudy (goldeverything):

ok simple

OpenStudy (anonymous):

just show me what to do

OpenStudy (goldeverything):

\[x +5=3x -5\]

OpenStudy (goldeverything):

your going to want to subtract 5 from each side so it would look like this

OpenStudy (goldeverything):

\[x =x3-10\]

OpenStudy (anonymous):

5

OpenStudy (anonymous):

x=5

OpenStudy (goldeverything):

no

OpenStudy (anonymous):

oh im wrong hold on

OpenStudy (goldeverything):

and then your going to want to drop both of the x's and then it is a addition problem

OpenStudy (goldeverything):

\[3+10=?\]

OpenStudy (goldeverything):

now what is the awnser

OpenStudy (anonymous):

i did everything you told me to do and i still got 5 for some reason

OpenStudy (goldeverything):

use the equation i just did

OpenStudy (anonymous):

13

OpenStudy (goldeverything):

correct

OpenStudy (anonymous):

im so dumb lmbo thank you

OpenStudy (goldeverything):

we all have are moments it's fine were only human